Monica is 4 years older her brother Kevin. Kevin is 9 years younger than Robert, the oldest child in the family. The sum of their ages is 37. How old are Robert, Monica, and Kevin
Found 3 solutions by ikleyn, josgarithmetic, MathTherapy:
Answer by ikleyn(52788)

.

    Robert + Kevin + Monica    = 37


      x    + (x-9) + ((x-9)+4) = 37


      x    +  x-9  +   x-5     = 37


          3x   -  14           = 37


          3x                   = 37 + 14 = 51


           x                             = 51/3 = 17.


ANSWER.  Robert is 17 years old;  Kevin is  17-9 = 8 years old;  Monica is 8+4 = 12 years old.

Solved.



Answer by josgarithmetic(39618)

Monica          r-9+4
Kevin           r-9
Robert          r
SUM              37

Answer by MathTherapy(10552)


Monica is 4 years older her brother Kevin. Kevin is 9 years younger than Robert, the oldest child in the family. The sum of their ages is 37. How old are Robert, Monica, and Kevin
Let Kevin's age be K
Then Monica's is: K + 4, and Robert's is: K + 9
We then get: K + K + 4 + K + 9 = 37
3K = 24
Kevin, or
Now you should be able to find Monica's and Robert's ages.